Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 574
Entry time: 16 Dec 2024, 09:44
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 16.12.2024 |
|
|
General Announcement:
- Next Wednesday it is the last XO Integration Meeting.
CTRL:
- Proceeding with Karabo update accordingly to the schedule.
DA:
- Nothing to report.
EEE:
- Issue with PLCs: on Friday the SA3 GATT Loop had issue; not fixed until now and loop will be brought back as soon as possible. Working is ongoing.
VAC:
- Nothing to report.
XPD:
- Proceeding with:
PES repair (today afternoon goes back to tunnel);
SA1 XFELO photon diode and grating was installed (electronics ADC card in the uTCA crate was not found in the crate and EEE was contacted – P. Gessler says a normal card will be installed in the coming
days).
DET:
- Nothing to report; tests to be performed after Karabo updates.
ITDM:
- SA1 control network switches updates were done (two switches not yet operational but Desy experts are working on it).
XRO:
- Progressing well with maintenance activities.
- Note to CTRL: today Karabo update in SA1, tomorrow in SA2 and Wednesday in SA3; Also EEE makes an update today in SA3.
LAS:
- Not present.
UNSYS:
- Nothing to report.
SPB:
- Nothing to report.
FXE:
- Nothing to report.
- LPD burned cable and plugs: new connectors were ordered and will be replaced.
MID:
- All fine.
HED:
- Nothing to report.
SCS:
- Nothing to report.
SQS:
- Not present.
SXP:
- Not present. |